# and Large Deviation Bounds

### v2

@@ 1. Introduction
Contributions. Our main contributions are built on several pillars. First, we analyze the strongly convex quadratic setting under i.i.d. isotropic Gaussian gradient noise. This case provides a clean and tractable foundation, allowing us to develop explicit characterizations before addressing the more general class of smooth, strongly convex functions subject to potentially biased noise. In the quadratic setting, we investigate how the risk-sensitive index $R(\theta)$ depends on the given algorithm parameters $(\alpha,\beta,\nu)$ and risk-aversion parameter $\theta$. By modeling the algorithmic class GMM as a linear time-invariant dynamical system and leveraging some existing results from robust control theory, we argue that $R(\theta)$ admits an explicit characterization through solving a discrete-time algebraic Riccati equation (DARE) that depends on the algorithm parameters. However, this is a matrix equation over $2d\times 2d$ matrices, and in general, it entails solving a $4d\times 4d$ generalized eigenvalue problem, which is computationally prohibitive for large $d$ requiring at least $\mathcal{O}(d^{3})$ operations. To this end, we establish a dimension-reduction result: for quadratic objectives subject to i.i.d. isotropic Gaussian noise with variance $\sigma^{2}$, we show that the DARE decomposes into $d$ independent $2\times 2$ Riccati equations. Each such equation can be solved in $\mathcal{O}(1)$ time, allowing the overall solution to be computed in $\mathcal{O}(d)$ operations in parallel. This yields a closed-form expression for the risk-sensitive index in terms of easily solvable $2\times 2$ DAREs. Furthermore, for gradient descent, we can solve the DAREs explicitly and provide an explicit expression for the risk-sensitive index. Building on our characterization of the risk-sensitive index, we then establish a fundamental trade-off between risk (as measured by this index) and convergence rate (quantifying solution speed in the absence of noise) in momentum methods. In particular, parameter choices that slow down convergence speed also reduce the associated risk and enlarge the range of $\theta$ values for which the risk remains finite, i.e., improved risk sensitivity and greater flexibility in $\theta$ comes at the expense of a reduction in the convergence rate. Specifically, the risk is finite if and only if $\theta>0$ satisfies $\sqrt{\theta}\,H_{\infty}<\sqrt{d},$ where $H_{\infty}$ denotes the $H_{\infty}$-norm of the dynamical system associated with the GMM iterations—explicitly characterized for quadratic objectives in [57]. Because the $H_{\infty}$-norm quantifies robustness to the worst-case deterministic gradient errors [57], this relation implies that greater robustness, i.e., smaller $H_{\infty}$-norm, permits a wider range of $\theta$ for which the risk is finite, thereby linking risk sensitivity (a probabilistic performance measure) to deterministic worst-case robustness in momentum-based methods. While connections between risk sensitivity and the $H_{\infty}$-norm exist in the context of control systems and controlled Markov processes [63, Sec. VI], [61, 64]; what these connections mean in the context of momentum-based optimization algorithms subject to gradient errors is first explored in our work to our knowledge. We then compare momentum methods—HB, NAG, GD, and TMM—across different parameter settings, focusing on their risk profiles, i.e., how their risk sensitivity varies with stepsize and momentum. We find that HB with standard parameters (proposed by Polyak [8]) converges faster than NAG (with standard parameters as in [7]) but incurs higher risk, whereas GD (with the common stepsize $\alpha=1/L$ where $L$ is the Lipschitz constant of the gradient) is slower yet exhibits a lower risk than both methods. These trade-offs trace a Pareto-optimal frontier over parameter choices, delivering the best attainable risk for a target convergence rate –we illustrate this frontier numerically by leveraging our dimension-reduction framework to compute the risk index.
@@ 2. Related Work
Compared to in-expectation guarantees, high-probability guarantees under biased noise are less developed. While in-expectation guarantees can be converted into high-probability guarantees using strategies such as longer runs of the algorithm combined with Chebyshev’s inequality, or parallel runs with selection of the best outcome [80, p. 243], [81], these strategies often produce loose complexity bounds with respect to the target probability and desired precision and can be impractical in streaming data settings [82], [56, Remark 13]. Existing high-probability results include random reshuffling for SGD [31, 30] and non-linear SGD variants with heavy-tailed noise where clipping, quantization, or normalization induce bias [83, 68].
@@
Next, we are going to discuss a related body of work from the risk perspective. Risk sensitivity, and more broadly, entropic risk, has been covered in the literature largely in the control theory context [59, 92, 60, 61]. These works consider centered, independent and identically distributed (i.i.d.) Gaussian noise. In particular, in [61], an exact analysis of $R(\theta)$ is provided in the linear–quadratic control setting. These results are not given in the context of optimization algorithms but rather for general linear dynamical systems. In our work, however, we interpret GMM as a linear dynamical system when the objective is a quadratic and show that their result translates directly into a formula for the risk-sensitive cost of GMM algorithms in this setting. However, computing the risk-sensitivity index via the approach of [61] requires integrating the system’s transfer function in the frequency domain, which—using the results of [64], [93, Remark 4.11], and [94, Lemma 3.4]—can be reduced to solving $2d\times 2d$ Riccati equations, entailing at least $\mathcal{O}(d^{3})$ computational complexity. By exploiting the special structure of the GMM, we show that it suffices to solve $2\times 2$ Riccati equations, which involve three unknowns—the entries of a symmetric $2\times 2$ matrix—and three corresponding equations; moreover, the approach in [61] applies only to quadratics and does not extend to strongly convex objectives. More recently, in [53] the authors introduced robustness measures related to asymptotic optimality under unbiased gradient errors, analyzing quadratic functions and, more generally, smooth strongly convex functions. In particular, the trade-offs between convergence rate and robustness to gradient errors for GD and NAG have been analyzed in [53]; however, in contrast to ours, the approach in [53] does not extend to biased gradient estimates or deterministic worst-case errors –the rate vs risk trade-off scheme introduced in [53] for first-order optimization methods is later extended in [55] to analyze first-order primal-dual methods for solving strongly convex-strongly concave saddle point problems subject to unbiased gradient errors with finite variance. In a different direction from [53], a more recent work [58] proposes an alternative robustness measure motivated by the entropic risk of the iterates and provides high-probability bounds for GMM in the unbiased sub-Gaussian gradient noise setting. However, their analysis does not accommodate biased gradient estimates, focuses solely on the suboptimality of the last iterate $x_{K}$ rather than on ergodic averages or cumulative suboptimality $S_{K}$, and does not establish large-deviation bounds. Compared to prior work, we treat (i) constant step sizes (not decaying), (ii) biased sub-Gaussian gradient errors (not just unbiased), and (iii) risk-sensitive metrics (RSI/LDP) and their link to $H_{\infty}$-norm in the context of momentum-based optimization algorithms.
@@ 3. Preliminaries
Note that in ˜3.1, the noise sequence can be biased; i.e., $\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{k+1}}|\mathcal{F}_{k}]$ =0=0 does not necessarily hold. For example, when $w_{k+1}=\tilde{\zeta}_{k+1}+\tilde{\delta}_{k}$ where $\tilde{\zeta}_{k+1}$ is i.i.d. sub-Gaussian, independent from $\mathcal{F}_{k}$ with variance proxy ${\tilde{\sigma}}^{2}$ and mean $\mathbb{E}[\tilde{\zeta}_{k+1}|\mathcal{F}_{k}]=0$ and $\tilde{\delta}_{k}$ is a bounded deterministic sequence with the bound $\|\tilde{\delta}_{k}\|^{2}\leq M^{2}$ for some $M\geq 0$; then, it is easy to see that ˜3.1 holds for $\sigma^{2}=\bar{c}_{0}(\tilde{\sigma}^{2}+M^{2})$, for some large enough constant $\bar{c}_{0}>0$, where we did not attempt to optimize the choice of $\sigma$. As a special case, if $w:=\{w_{k+1}\}_{k\geq 0}\in\ell_{2}(\mathbb{R}^{d})$—that is, if $\{w_{k+1}\}_{k\geq 0}$ is a deterministic square-summable sequence satisfying $\sum_{k\geq 0}\|w_{k}\|^{2}<\infty$—then such a finite $M$ exists and ˜3.1 is satisfied.
@@ 3.1. Risk-sensitive cost and the risk-sensitive index
The $H_{\infty}$-norm quantifies the worst-case deterministic noise amplification of the system in (3.8), i.e., it quantifies how the output behaves in response to the noise input $\{w_{k+1}\}_{k\geq 0}$. A smaller $H_{\infty}$ value indicates a more robust system, as the output is less sensitive to worst-case deterministic square-summable input noise; consequently, the cumulative suboptimality over the iterations, i.e., $\sum_{k\geq 0}\big(f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big)$, will also be smaller for the same noise level. For GMM, $H_{\infty}$-norm depends on the algorithm parameters $(\alpha,\beta,\nu)$ and the problem parameters $L$ and $\mu$, i.e., smoothness and convexity constants, respectively. For strongly convex quadratics, it is known that $H_{\infty}$-norm can be computed exactly. The following result is a direct consequence of [57, Thm. 4.1] together with [57, eqns. (60)–(61)].
@@
For general strongly convex functions that are not quadratic, exact analytical expressions are typically unavailable. Nevertheless, tight upper bounds have been derived in [57] for GD, NAG, and more generally for GMM, by verifying the feasibility of certain $4\times 4$ matrix inequalities. For completeness, we present these bounds in theorem˜A.1 in the appendix. In the context of stochastic dynamical systems driven by Gaussian noise, it is known that risk sensitivity is closely related to the $H_{\infty}$-norm: specifically, the largest value of $\theta>0$ for which the risk-sensitive index $R(\theta)$ remains finite depends on the system’s $H_{\infty}$-norm [106, 61]. To the best of our knowledge, however, such connections have not been explored in the setting of optimization algorithms driven by noise models such as those described in ˜3.1. To investigate this connection further and gain insight via precise analytical formulas, the next section focuses on the special case of strongly convex quadratic objectives subject to unbiased (centered) i.i.d. Gaussian noise. This serves as a foundation before extending the analysis to general smooth strongly convex functions under the broader noise model of ˜3.1.
@@ 4. Exact analysis of $R(\theta)$ for strongly convex quadratics
and, from (4.3), it can be seen that $z_{k}$ satisfies the identity (3.13). Therefore, based on eq.˜3.14, we can talk about the $H_{\infty}$-norm of the GMM system and the $H_{\infty}$ formula from theorem˜3.4 is applicable. The $H_{\infty}$-norm characterizes robustness in the hypothetical setting where the noise is fully deterministic. In ˜3.1, we consider a more general noise model; however, if the noise specified therein were further restricted to be deterministic and square-summable, the $H_{\infty}$-norm would quantify robustness with respect to the worst-case impact of such noise.
The GMM dynamics, as defined by (4.5)–(4.7), constitute a linear dynamical system identified by the system matrices $(A_{Q},B,T)$. In this context, in the special case when the noise sequence $\{w_{k+1}\}_{k\geq 0}$ consists of i.i.d. multivariate Gaussian vectors such that $w_{k+1}\sim\mathcal{N}(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\sigma^{2}}{d}{\mathsf{I}_{d}}}})$ for some $\sigma>0$, then the noise variance is $\sigma^{2}$ and ˜3.1 is satisfied. In this case, it is known that the risk-sensitive index $R(\theta)$ is finite if and only if $\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}}$ [61]; indeed, a particular form of a discrete-time algebraic Riccati equation (DARE), i.e.,
@@
admits a positive semi-definite solution $X\succeq 0$ if and only if $\gamma>H_{\infty}$ [112, Corollary 2.1, Prop. 3.7] –thus, $\sqrt{\theta}H_{\infty}<\sqrt{d}$ holds if and only if (4.8) has a positive semi-definite solution for the parameter choice $\gamma=\frac{1}{\sqrt{\theta/d}}$ (see [94, Lemma 3.4] and [61]). To compute a solution $X\in\mathbb{R}^{2d\times 2d}$ for the DARE above, a standard method involves tackling a $4d\times 4d$ generalized eigenvalue problem. More specifically, it necessitates the calculation of the generalized eigenvalues and eigenvectors of the following $4d\times 4d$ matrix pencil:

### v3

@@ 1. Introduction
Contributions. Our main contributions are built on several pillars. First, we analyze the strongly convex quadratic setting under i.i.d. isotropic Gaussian gradient noise. This case provides a clean and tractable foundation, allowing us to develop explicit characterizations before addressing the more general class of smooth, strongly convex functions subject to potentially biased noise. In the quadratic setting, we investigate how the risk-sensitive index $R(\theta)$ depends on the given algorithm parameters $(\alpha,\beta,\nu)$ and risk-aversion parameter $\theta$. By modeling the algorithmic class GMM as a linear time-invariant dynamical system and leveraging some existing results from robust control theory, we argue that $R(\theta)$ admits an explicit characterization through solving a discrete-time algebraic Riccati equation (DARE) that depends on the algorithm parameters. However, this is a matrix equation over $2d\times 2d$ matrices, and in general, it entails solving a $4d\times 4d$ generalized eigenvalue problem, which is computationally prohibitive for large $d$ requiring at least $\mathcal{O}(d^{3})$ operations. To this end, we establish a dimension-reduction result: for quadratic objectives subject to i.i.d. isotropic Gaussian noise with variance $\sigma^{2}$, we show that the DARE decomposes into $d$ independent $2\times 2$ Riccati equations. Each such equation can be solved in $\mathcal{O}(1)$ time, allowing the overall solution to be computed in $\mathcal{O}(d)$ operations in parallel. This yields a closed-form expression for the risk-sensitive index in terms of easily solvable $2\times 2$ DAREs. Furthermore, for gradient descent, we can solve the DAREs explicitly and provide an explicit expression for the risk-sensitive index. Building on our characterization of the risk-sensitive index, we then establish a fundamental trade-off between risk (as measured by this index) and convergence rate (quantifying solution speed in the absence of noise) in momentum methods. In particular, parameter choices that slow down convergence speed also reduce the associated risk and enlarge the range of $\theta$ values for which the risk remains finite, i.e., improved risk sensitivity and greater flexibility in $\theta$ comes at the expense of a reduction in the convergence rate. Specifically, the risk is finite if and only if $\theta>0$ satisfies $\sqrt{\theta}\,H_{\infty}<\sqrt{d},$ where $H_{\infty}$ denotes the $H_{\infty}$-norm of the dynamical system associated with the GMM iterations—explicitly characterized for quadratic objectives in [58]. Because the $H_{\infty}$-norm quantifies robustness to the worst-case deterministic gradient errors [58], this relation implies that greater robustness, i.e., smaller $H_{\infty}$-norm, permits a wider range of $\theta$ for which the risk is finite, thereby linking risk sensitivity (a probabilistic performance measure) to deterministic worst-case robustness in momentum-based methods. While connections between risk sensitivity and the $H_{\infty}$-norm exist in the context of control systems and controlled Markov processes [41, Sec. VI], [46, 66]; what these connections mean in the context of momentum-based optimization algorithms subject to gradient errors is first explored in our work to our knowledge. We then compare momentum methods—HB, NAG, GD, and TMM—across different parameter settings, focusing on their risk profiles, i.e., how their risk sensitivity varies with stepsize and momentum. We find that HB with standard parameters (proposed by Polyak [90]) converges faster than NAG (with standard parameters as in [85]) but incurs higher risk, whereas GD (with the common stepsize $\alpha=1/L$ where $L$ is the Lipschitz constant of the gradient) is slower yet exhibits a lower risk than both methods. These trade-offs trace a Pareto-optimal frontier over parameter choices, delivering the best attainable risk for a target convergence rate –we illustrate this frontier numerically by leveraging our dimension-reduction framework to compute the risk index.
@@ 2. Related Work
Compared to in-expectation guarantees, high-probability guarantees under biased noise are less developed. While in-expectation guarantees can be converted into high-probability guarantees using strategies such as longer runs of the algorithm combined with Chebyshev’s inequality, or parallel runs with selection of the best outcome [83, p. 243], [63], these strategies often produce loose complexity bounds with respect to the target probability and desired precision and can be impractical in streaming data settings [49], [74, Remark 13]. Existing high-probability results include random reshuffling for SGD [54, 114] and non-linear SGD variants with heavy-tailed noise where clipping, quantization, or normalization induce bias [23, 3].
@@
Next, we are going to discuss a related body of work from the risk perspective. Risk sensitivity, and more broadly, entropic risk, has been covered in the literature largely in the control theory context [68, 109, 108, 46]. These works consider centered, independent and identically distributed (i.i.d.) Gaussian noise. In particular, in [46], an exact analysis of $R(\theta)$ is provided in the linear–quadratic control setting. These results are not given in the context of optimization algorithms but rather for general linear dynamical systems. In our work, however, we interpret GMM as a linear dynamical system when the objective is a quadratic and show that their result translates directly into a formula for the risk-sensitive cost of GMM algorithms in this setting. However, computing the risk-sensitivity index via the approach of [46] requires integrating the system’s transfer function in the frequency domain, which—using the results of [66], [87, Remark 4.11], and [100, Lemma 3.4]—can be reduced to solving $2d\times 2d$ Riccati equations, entailing at least $\mathcal{O}(d^{3})$ computational complexity. By exploiting the special structure of the GMM, we show that it suffices to solve $2\times 2$ Riccati equations, which involve three unknowns—the entries of a symmetric $2\times 2$ matrix—and three corresponding equations; moreover, the approach in [46] applies only to quadratics and does not extend to strongly convex objectives. More recently, in [6] the authors introduced robustness measures related to asymptotic optimality under unbiased gradient errors, analyzing quadratic functions and, more generally, smooth strongly convex functions. In particular, the trade-offs between convergence rate and robustness to gradient errors for GD and NAG have been analyzed in [6]; however, in contrast to ours, the approach in [6] does not extend to biased gradient estimates or deterministic worst-case errors –the rate vs risk trade-off scheme introduced in [6] for first-order optimization methods is later extended in [116] to analyze first-order primal-dual methods for solving strongly convex-strongly concave saddle point problems subject to unbiased gradient errors with finite variance. In a different direction from [6], a more recent work [21] proposes an alternative robustness measure motivated by the entropic risk of the iterates and provides high-probability bounds for GMM in the unbiased sub-Gaussian gradient noise setting. However, their analysis does not accommodate biased gradient estimates, focuses solely on the suboptimality of the last iterate $x_{K}$ rather than on ergodic averages or cumulative suboptimality $S_{K}$, and does not establish large-deviation bounds. Compared to prior work, we treat (i) constant step sizes (not decaying), (ii) biased sub-Gaussian gradient errors (not just unbiased), and (iii) risk-sensitive metrics (RSI/LDP) and their link to $H_{\infty}$-norm in the context of momentum-based optimization algorithms.
@@ 3. Preliminaries
Note that in assumption˜3.1, the noise sequence can be biased; i.e., $\mathbb{E}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w_{k+1}}|\mathcal{F}_{k}]$ =0=0 does not necessarily hold. For example, when $w_{k+1}=\tilde{\zeta}_{k+1}+\tilde{\delta}_{k}$ where $\tilde{\zeta}_{k+1}$ is i.i.d. sub-Gaussian, independent from $\mathcal{F}_{k}$ with variance proxy ${\tilde{\sigma}}^{2}$ and mean $\mathbb{E}[\tilde{\zeta}_{k+1}|\mathcal{F}_{k}]=0$ and $\tilde{\delta}_{k}$ is a bounded deterministic sequence with the bound $\|\tilde{\delta}_{k}\|^{2}\leq M^{2}$ for some $M\geq 0$; then, it is easy to see that assumption˜3.1 holds for $\sigma^{2}=\bar{c}_{0}(\tilde{\sigma}^{2}+M^{2})$, for some large enough constant $\bar{c}_{0}>0$, where we did not attempt to optimize the choice of $\sigma$. As a special case, if $w:=\{w_{k+1}\}_{k\geq 0}\in\ell_{2}(\mathbb{R}^{d})$—that is, if $\{w_{k+1}\}_{k\geq 0}$ is a deterministic square-summable sequence satisfying $\sum_{k\geq 0}\|w_{k}\|^{2}<\infty$—then such a finite $M$ exists and assumption˜3.1 is satisfied.
@@ 3.1. Risk-sensitive cost and the risk-sensitive index
The $H_{\infty}$-norm quantifies the worst-case deterministic noise amplification of the system in (3.8), i.e., it quantifies how the output behaves in response to the noise input $\{w_{k+1}\}_{k\geq 0}$. A smaller $H_{\infty}$ value indicates a more robust system, as the output is less sensitive to worst-case deterministic square-summable input noise; consequently, the cumulative suboptimality over the iterations, i.e., $\sum_{k\geq 0}\big(f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big)$, will also be smaller for the same noise level. For GMM, $H_{\infty}$-norm depends on the algorithm parameters $(\alpha,\beta,\nu)$ and the problem parameters $L$ and $\mu$, i.e., smoothness and convexity constants, respectively. For strongly convex quadratics, it is known that $H_{\infty}$-norm can be computed exactly. The following result is a direct consequence of [58, Thm. 4.1] together with [58, eqns. (60)–(61)].
@@
For general strongly convex functions that are not quadratic, exact analytical expressions are typically unavailable. Nevertheless, tight upper bounds have been derived in [58] for GD, NAG, and more generally for GMM, by verifying the feasibility of certain $4\times 4$ matrix inequalities. For completeness, we present these bounds in theorem˜A.1 in the appendix. In the context of stochastic dynamical systems driven by Gaussian noise, it is known that risk sensitivity is closely related to the $H_{\infty}$-norm: specifically, the largest value of $\theta>0$ for which the risk-sensitive index $R(\theta)$ remains finite depends on the system’s $H_{\infty}$-norm [42, 46]. To the best of our knowledge, however, such connections have not been explored in the setting of optimization algorithms driven by noise models such as those described in assumption˜3.1. To investigate this connection further and gain insight via precise analytical formulas, the next section focuses on the special case of strongly convex quadratic objectives subject to unbiased (centered) i.i.d. Gaussian noise. This serves as a foundation before extending the analysis to general smooth strongly convex functions under the broader noise model of assumption˜3.1.
@@ 4. Exact analysis of $R(\theta)$ for strongly convex quadratics
and, from (4.3), it can be seen that $z_{k}$ satisfies the identity (3.13). Therefore, based on eq.˜3.14, we can talk about the $H_{\infty}$-norm of the GMM system and the $H_{\infty}$ formula from theorem˜3.4 is applicable. The $H_{\infty}$-norm characterizes robustness in the hypothetical setting where the noise is fully deterministic. In assumption˜3.1, we consider a more general noise model; however, if the noise specified therein were further restricted to be deterministic and square-summable, the $H_{\infty}$-norm would quantify robustness with respect to the worst-case impact of such noise.
The GMM dynamics, as defined by (4.5)–(4.7), constitute a linear dynamical system identified by the system matrices $(A_{Q},B,T)$. In this context, in the special case when the noise sequence $\{w_{k+1}\}_{k\geq 0}$ consists of i.i.d. multivariate Gaussian vectors such that $w_{k+1}\sim\mathcal{N}(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\sigma^{2}}{d}{\mathsf{I}_{d}}}})$ for some $\sigma>0$, then the noise variance is $\sigma^{2}$ and assumption˜3.1 is satisfied. In this case, it is known that the risk-sensitive index $R(\theta)$ is finite if and only if $\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}}$ [46]; indeed, a particular form of a discrete-time algebraic Riccati equation (DARE), i.e.,
@@
admits a positive semi-definite solution $X\succeq 0$ if and only if $\gamma>H_{\infty}$ [62, Corollary 2.1, Prop. 3.7] –thus, $\sqrt{\theta}H_{\infty}<\sqrt{d}$ holds if and only if (4.8) has a positive semi-definite solution for the parameter choice $\gamma=\frac{1}{\sqrt{\theta/d}}$ (see [100, Lemma 3.4] and [46]). To compute a solution $X\in\mathbb{R}^{2d\times 2d}$ for the DARE above, a standard method involves tackling a $4d\times 4d$ generalized eigenvalue problem. More specifically, it necessitates the calculation of the generalized eigenvalues and eigenvectors of the following $4d\times 4d$ matrix pencil: